OpenStudy (anonymous):

@satellite73 A bag contains red and blue marbles. The probability of drawing a red marble is 40 percent. What is the probability of drawing a blue marble? A. 20% B. 40% C. 60% D. 80%

OpenStudy (anonymous):

i think it is A

OpenStudy (anonymous):

what is \(100\%-40\%\)

OpenStudy (anonymous):

no, it is not A

OpenStudy (anonymous):

100-40=60

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Another way to think about this is that if you had 10 marbles in the bag, 4 out of the 10 would be red. Therefore, 6 out of 10 would be blue. Meaning that you would have a 60% chance.

OpenStudy (anonymous):

Right? @satellite73

OpenStudy (anonymous):

can u help me with this question please @satellite73

OpenStudy (anonymous):

Determine the sum of the given probabilities. Are the events complementary? P(gold) =4/11; P(silver) = 7/11 A. 1 The events are complementary. B. 1 The events are not complementary. C.1/2 The events are complementary. D.1/2The events are not complementary.

OpenStudy (anonymous):

@satellite73 help please

OpenStudy (anonymous):

@Zarkon help

OpenStudy (anonymous):

complimentary means either one happens or the other

OpenStudy (anonymous):

i.e. one of them is sure to happen like either it rains or it does not you roll an even number or you roll an odd number you flip head or you flip tails those are examples of complimentary events

OpenStudy (anonymous):

so A. am i right @satellite73 or @Zarkon

OpenStudy (anonymous):

@sourwing @SolomonZelman @satellite73 @myininaya

OpenStudy (anonymous):

please help